OpenStudy (anonymous):

anyone in algebra 1b?quadric functions and equations unit test

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) Most prospective tutors and I would be more inclined to answer you quickly if you'd please post your homework problem immediately. then anyone willing to help you will help you.

OpenStudy (anonymous):

ok i have 2 questions

OpenStudy (anonymous):

A model rocket is launched from a roof into a large field.The apth of the rocket can be modeled by the equation y=-0.04x^2 + 8.3x + 4.3,where x is the horizontal distance in meters,from the starting point on the roof and y is the height in meters,of the rocket above the ground.How far horizontally from the starting point will the rocket land?Round your answer to the nearest hundredth meter. Options are A. 208.2 m B 416.03 m C.0.52 m D.208.19 m

OpenStudy (anonymous):

Landon is standing in a hole that is 6.5 ft deep. He throws a rock, and it goes up into the air, out of the hole, and then lands on the ground above. The path of the rock can be modeled by the equation y = –0.005x² + 0.45x – 6.5, where x is the horizontal distance of the rock, in meters, from Landon, and y is the height, in meters, of the rock above the ground. How far horizontally from Landon will the rock land? 18.07 ft 35.96 ft 9.04 ft 71.93 ft

jigglypuff314 (jigglypuff314):

where the rocket will land would be when height = y = 0 so solve for x when 0 = -0.04x^2 + 8.3x + 4.3

OpenStudy (anonymous):

How do i solve for x?im really bad at this

jigglypuff314 (jigglypuff314):

I suppose you would use the quadratic formula, have you learned that? :)

OpenStudy (anonymous):

No I get it mixed up/:

jigglypuff314 (jigglypuff314):

\[\large quadratic~formula = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]when in form ax^2 + bx + c = 0 so plug in a = -0.04 b = 8.3 and c = 4.3

OpenStudy (anonymous):

im still confused what do i do after i plug it in

jigglypuff314 (jigglypuff314):

\[\frac{ -8.3 + \sqrt{8.3^2 - (4*-0.04 * 4.3)} }{ 2* -0.04 }\]

jigglypuff314 (jigglypuff314):

I'm guessing you're allowed to use a calculator?

OpenStudy (anonymous):

okay I would solve this but I dont have my calculator

jigglypuff314 (jigglypuff314):

it would be the positive one

OpenStudy (anonymous):

Oh thank you so much

OpenStudy (anonymous):

so its d then?

jigglypuff314 (jigglypuff314):

yep :)

jigglypuff314 (jigglypuff314):

(now to copy paste so I don't have to scroll XD) Landon is standing in a hole that is 6.5 ft deep. He throws a rock, and it goes up into the air, out of the hole, and then lands on the ground above. The path of the rock can be modeled by the equation y = –0.005x² + 0.45x – 6.5, where x is the horizontal distance of the rock, in meters, from Landon, and y is the height, in meters, of the rock above the ground. How far horizontally from Landon will the rock land? 18.07 ft 35.96 ft 9.04 ft 71.93 ft

OpenStudy (anonymous):

its just 0.05x^2 with 2 zeros(:

OpenStudy (anonymous):

negative

jigglypuff314 (jigglypuff314):

so 0.05 not 0.005 ?

OpenStudy (anonymous):

yeah -0.05^2

OpenStudy (anonymous):

-0.05x^2*

jigglypuff314 (jigglypuff314):

are you sure??? :/ the -0.005x^2 had an answer -0.05x^2 doesn't have an answer...

OpenStudy (anonymous):

Yeah im sure whats the answer though?Is it any of the choices

jigglypuff314 (jigglypuff314):

ah yes it would be -0.005x^2 you would use the same idea and use logic to know that out of the two answers you could get, the right one would be when the line comes back down and hits the ground... http://www.wolframalpha.com/input/?i=0+%3D+%E2%80%930.005x%C2%B2+%2B+0.45x+%E2%80%93+6.5

OpenStudy (anonymous):

So its D then

jigglypuff314 (jigglypuff314):

yes :) but do you understand the idea behind plugging into the quadratic formula?

OpenStudy (anonymous):

Yup I just didn't remember the formula thanks alot though

jigglypuff314 (jigglypuff314):

glad I could help :)